Find the electric field at the center of the square given that:
Q1 = 3.4x10-13C
Q2 = Q3 = 1.5x10-16 C
Q4 = 4.6x10-12C
ℓ = .05m
1
Expert's answer
2020-07-10T10:26:13-0400

"E=\frac{k}{a^2}(Q_1+Q_2+Q_3+Q_4)\\E= \frac{9\times10^9}{(\frac{0.05}{\sqrt2})^2}(3.4\times10^{-13}+1.5\times10^{-16}+1.5\times10^{-16}+4.6\times10^{-12})"


"E=72\times10^2\times10^9\times4.94\times10^{-12}=35.56 N"